The sum of two numbers is 5 times their difference. If one exceeds the other by 7, what are the numbers?
Answer by mananth(16946)

let one number be x
the other number will be x+7
..
x+x+7 = 5(x+7-x)
2x+7 = 35
2x=28
x=14
the other number will be x+7 = 14+7 = 21